OpenStudy (anonymous):

which of the following is the solution to 4x^2-9=0

OpenStudy (anonymous):

-1/2, 3/2 1/2, -3/2 -3/2, 3/2 -3/2, 2

OpenStudy (michele_laino):

hint: \[\large 4{x^2} - 9 = \left( {2x - 3} \right)\left( {2x + 3} \right)\]